OpenStudy (anonymous):

what is the slope of (-5,-1) and (5,11)?

OpenStudy (anonymous):

6/5

OpenStudy (anonymous):

m=y2-y1/x2-x1 m=(11)-(-1)/(5)-(-5) m=12/10 (divide by 2:common multiple) m=6/5 Therefore the slope of (-5,-1) and (5,11) is 6/5